OpenStudy (anonymous):

If the margin of error in an estimate for the mean weight of a shipment is + or -2 pounds at a confidence level of 95 percent, what will be the margin of error at a confidence level of 98 percent? Be sure to show how you arrived at your answer.

OpenStudy (perl):

so we are given that margin of error is 2

OpenStudy (anonymous):

yes

OpenStudy (perl):

$$ \Large{ \text{confidence interval for sample mean is: } \\ \bar x \pm z^* \frac{\sigma }{\sqrt n} \\ Margin~ of~ error = z_{.05/2} \cdot \frac{\sigma }{\sqrt n} \\ \therefore \\ if ~ ~2 = 1.96 \cdot \frac{\sigma }{\sqrt n} \\ then ~~ \frac {2}{1.96 } = \frac{\sigma }{\sqrt n} \\ \therefore ME = z_{\frac{.02}{2}} \cdot \frac{\sigma }{\sqrt n} = z_{\frac{.02}{2}} \cdot \frac {2}{1.96 } } $$

OpenStudy (anonymous):

ok! what is the final answer though i am not sure how to compute it

OpenStudy (perl):

so you just need the critical score for 98% , two tailed

OpenStudy (perl):

$$\Large{ z_{\alpha/2}\\~~\\ z_{.10/2} = 1.645\\ z_{.05/2} = 1.960\\ z_{.02/2} = 2.326 } $$

OpenStudy (perl):